A) How many lone pairs are on the central atom in BCl3?

Express your answer numerically as an integer.

B)How many lone pairs are on the central atom of

BrF3?

Express your answer numerically as an integer.
Answer #1
Concepts and reason

Lone pair of electrons for an atom is defined as the non-bonding or unshared electrons in the outermost orbit of the atom.

Generally, these lone pair of electrons is shown in Lewis structure model of molecules.

In Lewis structures, non-bonding pair is located the central atom.

Fundamentals

Lewis structure: Lewis structure is a structure which shows all the bonding as well as non-bonding electrons present in it.

Example:

In this Lewis structure, the central atom is X, which has one lone pair in it and three bond pairs with 3 Y atoms.

Valence electrons

(central atom)

Number of bonds

(sigma bonds)

Number of lone pairs

P

Q

PQ2\frac{{{\rm{P - Q}}}}{{\rm{2}}}

(A)

The given molecule is BCl3{\rm{BC}}{{\rm{l}}_{\rm{3}}} .

In BCl3{\rm{BC}}{{\rm{l}}_{\rm{3}}} molecule, the central atom is boron (B).

The atomic number for boron atom is 5, and the valence electron for boron atom is 3.

The remaining chlorine atoms form sigma bonds with boron atom, and therefore, the total number of sigma bond present on the central atom is 3.

The number of lone pairs present on the central metal atom is calculated below.

Numberoflonepairs=(valenceelectrons)(numberofsigmabonds)2=(3)(3)2=0.\begin{array}{l}\\{\rm{Number}}\,{\rm{of}}\,{\rm{lone}}\,{\rm{pairs}}\,{\rm{ = }}\,\,\frac{{\left( {{\rm{valence}}\,{\rm{electrons}}} \right){\rm{ - }}\left( {{\rm{number}}\,{\rm{of}}\,{\rm{sigma}}\,{\rm{bonds}}} \right)}}{{\rm{2}}}\\\\\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm{ = }}\,\,\frac{{\left( {\rm{3}} \right){\rm{ - }}\left( {\rm{3}} \right)}}{{\rm{2}}}\\\\\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm{ = }}\,\,{\rm{0}}{\rm{.}}\\\end{array}

Lewis structure of the given molecule BCl3{\rm{BC}}{{\rm{l}}_{\rm{3}}} is shown below.

Therefore, the central boron atom has zero lone pair of electrons.

(B)

The given molecule is BrF3{\rm{Br}}{{\rm{F}}_{\rm{3}}} .

In BrF3{\rm{Br}}{{\rm{F}}_{\rm{3}}} molecule, the central atom is bromine (Br).

The atomic number for bromine atom is 35, and the valence electron for bromine atom is 7.

The remaining fluorine atoms form sigma bonds with bromine atom, and therefore, the total number of sigma bond present on the central atom is 3.

The number of lone pairs present on the central atom is calculated below.

Numberoflonepairs=(valenceelectrons)(numberofsigmabnds)2=(7)(3)2=42=2.\begin{array}{l}\\{\rm{Number}}\,{\rm{of}}\,{\rm{lone}}\,{\rm{pairs}}\,{\rm{ = }}\,\frac{{\left( {{\rm{valence}}\,{\rm{electrons}}} \right){\rm{ - }}\left( {{\rm{number}}\,{\rm{of}}\,{\rm{sigma}}\,{\rm{bnds}}} \right)}}{{\rm{2}}}\\\\\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm{ = }}\,\frac{{\left( {\rm{7}} \right){\rm{ - }}\left( {\rm{3}} \right)}}{{\rm{2}}}\\\\\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm{ = }}\,\frac{{\rm{4}}}{{\rm{2}}}\\\\\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm{ = }}\,2.\\\end{array}

Lewis structure of the given molecule BrF3{\rm{Br}}{{\rm{F}}_{\rm{3}}} is shown below.

:i:

Therefore, the central bromine atom has two lone pair of electrons.

Ans: Part A

The number of lone pair electron present on central atom in BCl3{\rm{BC}}{{\rm{l}}_{\rm{3}}} is 0

Part B

The number of lone pair electron present on the central atom in BrF3{\rm{Br}}{{\rm{F}}_{\rm{3}}} is 2.